# WISE: Rethinking the Knowledge Memory for Lifelong Model Editing of Large Language Models

### v1

@@ Abstract
Large language models (LLMs) need knowledge updates to meet the ever-growing world facts and correct the hallucinated responses, facilitating the methods of lifelong model editing. Where the updated knowledge resides in memories is a fundamental question for model editing. In this paper, we find that editing either long-term memory (direct model parameters) or working memory (non-parametric knowledge of neural network activations/representations by retrieval) will result in an impossible triangle—reliability, generalization, and locality can not be realized together in the lifelong editing settings. For long-term memory, directly editing the parameters will cause conflicts with irrelevant pretrained knowledge or previous edits (poor reliability and locality). For working memory, retrieval-based activations can hardly make the model understand the edits and generalize (poor generalization). Therefore, we propose WISE to bridge the gap between memories. In WISE, we design a dual parametric memory scheme, which consists of the main memory for the pretrained knowledge and a side memory for the edited knowledge. We only edit the knowledge in the side memory and train a router to decide which memory to go through when given a query. For continual editing, we devise a knowledge-sharding mechanism where different sets of edits reside in distinct subspaces of parameters, and are subsequently merged into a shared memory without conflicts. Extensive experiments show that WISE can outperform previous model editing methods and overcome the impossible triangle under lifelong model editing of question answering, hallucination, and out-of-distribution settings across trending LLM architectures, e.g., GPT, LLaMA, and Mistral [^1] Code will be released at https://github.com/zjunlp/EasyEdit..
@@ 1 Introduction
Large language models (LLMs) show emergent intelligence when scaling the number of parameters, computes, and data [1, 2, 3, 4], which reveals the sparks of artificial general intelligence [5]. However, when deployed, LLMs still make mistakes [6], generating responses with hallucinations [7], bias [8], and factual decays [9]. On the other hand, the world’s knowledge is ever-growing, so the up-to-date knowledge is usually different from the one during LLMs’ pretraining [10]. Many such errors and emerging facts will arise sequentially in deployment, some of which have to be addressed timely and efficiently without waiting for retraining or finetuning [11, 12]. Also, retraining or finetuning is often too computationally expensive [13, 10], which is not sustainable for lifelong growing knowledge. Therefore, lifelong model editing [10] was proposed to remedy the continual knowledge updates and injections for LLMs in a cheap and timely manner.
An effective lifelong model editing approach should satisfy the following properties [14, 15, 11]: i) reliability, the model can remember both current and previous edits after sequential editing; ii) locality, model editing will not influence inherent pretrained knowledge which is irrelevant to the edited knowledge; iii) generalization, the model is not just merely memorizing the query-target pairs;
@@ 2.3.1 Side Memory Design
Side memory in FFN’s value matrix. Each layer in a Transformer contains a multi-head self-attention (MHA) mechanism and a feed-forward network (FFN), where the FFN constitutes two-thirds of the model parameters [31]. The question of how Transformers retrieve and utilize stored knowledge remains unresolved [16, 32], yet past works [29, 31] have demonstrated that editing the weights of the FFN is consistently more effective for LLMs. The FFN typically consists of key-value linear matrices: $\mathbf{W}_{k},\mathbf{W}_{v}$, i.e., two multi-layer perceptron (MLP) layers. For the output of attention feature $\mathbf{f}$, the computation of the feed-forward network, omitting the bias terms, can be represented as:
@@
Locating side memory’s FFN layer. Transformer LLMs have been widely demonstrated to encode “lower-level” information (e.g., parts of speech) in earlier layers while processing more advanced linguistic phenomena like anaphora and coreference in later layers [33, 34, 35]. Representations in later hidden layers propagate through residual connections without drastic changes [36, 16], enabling effective early exit in LLMs [37, 38]. Therefore, to minimize the side effects of editing and adjust advanced linguistic phenomena, we target mid-to-late layers (e.g. 27) for side memory. Further analysis of layer selection is provided in Section 5.
Routing between side memories and main memory. Similar to the retrieval-based methods [10, 30], during inference, it is needed to decide whether the main memory or the side memory is used. If a given query is within the scope of previous edits, the side memory is used; otherwise, the main memory. Inspired by [11], we introduce a routing activation indicator, given an input $\mathbf{x}$, it is formulated:
@@ 2.3.2 Knowledge Sharding and Merging
Knowledge in random memory subspaces. We edit the side memory $\mathbf{W}_{v^{\prime}}$. We divide $n$ edits into $k$ shards, copy the side memory for $k$ times, and generate $k$ random gradient mask with mask ratio $\rho$ for each copy of side memory. A random gradient mask $\mathbf{M}_{i}\in\{0,1\}^{|\mathbf{W}_{v^{\prime}}|},i\in[k]$ is a binary mask whose proportion of $1$ is $\rho$ [42]. For edit shard $i,i\in[k]$, we edit the knowledge into the subspace $\mathbf{M}_{i}$ as follows:
@@
Knowledge merging. We merge the $k$ subspace pieces of side memory into one. Because we randomly generate the subspace masks, different random masks will have some overlapping elements and some disjoint elements, following the theorem below:
@@
Routing and retrieving among several side memories. One single side memory has its limited knowledge capacity [41]. For the lifelong editing stream, we can produce several side memories and retrieve them via activation score routing. We compute different activation indicator scores of side memories and retrieve the top-1 during inference. This design is named WISE-Retrieve, which enables a more challenging lifelong editing scenario. For WISE with only one side memory, it is notated as WISE-Merge. For most of the experiments, we use WISE-Merge by default, and we compare WISE-Retrieve in Table 6 and Figure 6.
@@ 3.1 Experimental Settings and Evaluation Metrics
In the experiments, we compare the performance of different baselines and WISE in sequentially editing LLM models hundreds to thousands of times.
@@
Datasets and Models. We choose trending autoregressive LLM models LLaMA-2-7B [13], Mistral-7B [50], and GPT-J-6B [51, 52] for evaluation. The dataset details are in Table 3. Following [10], we evaluate WISE on the closed-book question-answering (QA) dataset ZsRE [44], and also evaluate its ability to correct Hallucination in SelfCheckGPT [46]. The Temporal dataset [48] is employed to test the out-of-distribution (OOD) generalization of editing. Since Temporal comprises emerging entities post-2019, we avoid using the latest LLMs in OOD experiments. Instead, we follow the original literature of the Temporal dataset [48] and adopt GPT-J-6B as the base model, which is pretrained on the Pile [49] with a cutoff in 2020. Implementation details and editing examples for each dataset and can be found in Appendix A.
Baselines. The baselines include methods of continual learning and model editing. We compare WISE against direct fine-tuning FT-L with an additional KL divergence loss [16], and continual learning fine-tuning based on Elastic Weight Consolidation (FT-EWC) [18]. We also compare WISE to other model editors, including 1) GPT-style editors based on causal tracing: ROME [16], MEMIT [17], and MEMIT-MASS (a batch-editing version of MEMIT); 2) hypernetwork-based editors: MEND [29]; and 3) the latest memory-based editors: DEFER (inspired by SERAC [30] for inference routing) and GRACE [10]. Details on all comparisons are found in Appendix A.2.
Metrics. Each edit example includes an edit descriptor (i.e., query) $\mathbf{x}_{e}$, its paraphrase prompts $\mathbf{x}_{e^{\prime}}$ (if available) for testing generalization, and an unrelated statement $\mathbf{x}_{\text{loc}}$ for testing locality. For the editing dataset $\mathcal{D}_{\text{edit}}=\{(\mathcal{X}_{e},\mathcal{Y}_{e})\}$ with $T$ edits, we evaluate the final post-edit model $f_{\Theta_{T}}$ after the $T$-th edit example $(\mathbf{x}_{T},\mathbf{y}_{T})$. We evaluate the model editor’s reliability and generalization using the metrics Rel. (a.k.a Edit Success Rate [10]) and Gen. (Generalization Success Rate [53]), while Loc. (Localization Success Rate [53]), defined as the post-edit model should not change the output of the irrelevant examples $\mathbf{x}_{\text{loc}}$, assesses specificity. We report these metrics and their mean scores, which are formally defined as:
@@ 3.2 Main Results
Competitive Performance of WISE. The competitive performance of WISE is evident in Table 2 and 4, which compare its results with eight baselines on the QA (ZsRE) and Hallucination (SelfCheckGPT) settings. In general, we observe the followings: ❶ WISE outperforms existing methods on multiple tasks after long editing sequences; ❷ direct editing of long-term memory (ROME, MEMIT, etc.) creates conflicts with prior pretraining knowledge, resulting in poor locality; and ❸ retrieving working memory and modifying activations (GRACE, DEFER, etc) struggle to generalize to diverse queries.
@@
Out-of-Distribution Evaluation. Ideally, model editing needs to generalize distributionally from formulaic editing examples to natural texts [48], where the distributional shift involves complexity rather than conventional domain shift [54]. Following [48], we evaluate the OOD generalization of editing methods on emerging entities using the temporal updating dataset, Temporal. Editing examples and evaluation metrics are provided in Appendix A.1. As shown in Table 5, WISE effectively handles out-of-distribution generalization tasks (achieving the best OOD Gen. and overall performance). DEFER delivers mediocre performance on OOD Gen. due to the limited capacity of the auxiliary model[14]. During the fine-tuning phase, GRACE and MEMIT focus on the representation $v*$ of a single input token after $\mathbf{W}_{v}$ (GRACE: last token, MEMIT: last subject token). However, regarding $v*$ the editing carrier encounters two problems: 1) the training objective is not aligned with the pretraining phase, and 2) the single representation limits the search scope of gradient descent, making it difficult to handle complex OOD generalization. WISE, on the other hand, avoids the above challenges.
@@ 3.3 Further Analysis
Visualization of WISE’s Routing Activation. To demonstrate the effectiveness of memory routing, we record the activation values $\Delta_{\text{act}}(\mathbf{x})$ of 1000 (QA, ZsRE)/600 (Halluc.) queries during the inference stage via knowledge merging into a single side memory. As shown in Figure 3, the purple horizontal line represents the activation threshold $\epsilon$ recorded during the editing phase. Almost all unrelated queries show low activations with values less than 10 in ZsRE and less than 20 in Halluc.; meanwhile, WISE accurately routes the editing prompt and unseen paraphrases into the side memory. This ensures editing locality and prevents excessive shifts from the pre-training distribution during lifelong editing.
@@
Localization Analysis of WISE’s Side Memory. To validate the benefits of editing mid-to-late layers, we select decoder layers from early, intermediate, mid-to-late, and late stages. As shown in Figure 5, the ablation results reveal that editing critical layers like the early and final layers (0, 1, 31) is ineffective, even resulting in a very low Loc. value of 0.096, which indicates a failure to recognize the editing scope. This may occur because the early layers represent fundamental grammatical information, and the final layer directly controls the decoding procedure, leading to poor editing of advanced language functions. Editing in the intermediate layers is suboptimal but still shows a markable improvement compared to early layers, possibly because intermediate layers start to integrate basic grammatical information with more complex semantic data. Notably, the mid-to-late layers demonstrate exceptional editing performance; for instance, selecting layer 26 results in an 80% success rate and generalization while maintaining 100% locality. This empirically supports our claim in Section 2.3.1 that the redundant mid-to-late layers [37] are ideal side memory layers and confirms the hierarchical nature of information processing in Transformer LLMs [55, 56].
Analysis of $\rho$ and $k$ for WISE. We analyze the important hyperparameters of WISE: the mask ratio $\rho$ and the number of subspaces $k$ in Figure 5. On the left figure, for $k=2$, the best $\rho$ is 0.2, satisfying $k*\rho=0.4<1$, which implies the effectiveness of our subspace design that higher knowledge density will cause better generalization. When scaling $k$, we observe an increasing demand of $\rho$. From Theorem 2.1, the probability of subspace overlap is $\rho^{k}$, and we hypothesize that this overlap is important as an anchor for model merging. Interestingly, from the right figure, it can be observed that the optimal cases always have the $\rho^{k}$ closest to 0.03. This shows an inherent tradeoff between merge anchor and merge conflicts, and the subspace overlaps around 0.03 are optimal for the best performances.
@@
Scale Up to 3K of Edits. We scale the number of continual edits to 3K in Table 6. We compare WISE-Merge, keeping one side memory by multi-time merging, and WISE-Retrieve, keeping several side memories by routing and retrieving among different side memories. For WISE-Retrieve, we show an upper bound “oracle”, which always identifies the correct routing path. We observe that the WISE series maintains high scalability, consistently outperforming the strongest baselines including MEMIT-MASS and GRACE. WISE-Retrieve based on top-1 activation retrieval demonstrates the best results in 3K edits, showing the effectiveness of well-organized memory subspaces and routing strategies during editing. We note that the “oracle” exhibits marginal performance decline when scaling the edits from 2K to 3K, yet it demonstrates remarkable performance across all metrics. This underscores the potential of WISE to handle extremely long continual edits, contingent upon substantial improvement in the retrieval of side memories. Additionally, an appropriate replay of edits can further improve retrieval accuracy, as detailed in Appendix B.3.
@@
Inference Time Analysis of WISE. Figure 6 shows the inference time of a single instance for LLaMA after $t\in[0,3000]$ editing steps, measured across 10 trials of each setting. Consistent with our expectations, we find that WISE-Merge incurs a constant inference delay (about 3%) as the editing stream expands. WISE-Retrieve, due to the introduction of retrieval routing, shows an increase in inference time as the number of edits increases, with a time cost increment of about 7% after 3K edits. Knowledge merging ensures that WISE-Merge only brings constant additional costs (0.64% extra parameters and 4% extra GPU VRAM, as detailed in Appendix B.6), contrasting with past memory-based works that continuously demand more available memory [10, 30].

### v2

@@ Abstract
Large language models (LLMs) need knowledge updates to meet the ever-growing world facts and correct the hallucinated responses, facilitating the methods of lifelong model editing. Where the updated knowledge resides in memories is a fundamental question for model editing. In this paper, we find that editing either long-term memory (direct model parameters) or working memory (non-parametric knowledge of neural network activations/representations by retrieval) will result in an impossible triangle—reliability, generalization, and locality can not be realized together in the lifelong editing settings. For long-term memory, directly editing the parameters will cause conflicts with irrelevant pretrained knowledge or previous edits (poor reliability and locality). For working memory, retrieval-based activations can hardly make the model understand the edits and generalize (poor generalization). Therefore, we propose WISE to bridge the gap between memories. In WISE, we design a dual parametric memory scheme, which consists of the main memory for the pretrained knowledge and a side memory for the edited knowledge. We only edit the knowledge in the side memory and train a router to decide which memory to go through when given a query. For continual editing, we devise a knowledge-sharding mechanism where different sets of edits reside in distinct subspaces of parameters, and are subsequently merged into a shared memory without conflicts. Extensive experiments show that WISE can outperform previous model editing methods and overcome the impossible triangle under lifelong model editing of question answering, hallucination, and out-of-distribution settings across trending LLM architectures, e.g., GPT, LLaMA, and Mistral [^1] Code is available at https://github.com/zjunlp/EasyEdit..
@@ 1 Introduction
Large language models (LLMs) show emergent intelligence when scaling the number of parameters and data [1, 2, 3, 4], which reveals the sparks of artificial general intelligence [5]. However, when deployed, LLMs still make mistakes [6], generating responses with hallucinations [7], bias [8], and factual decays [9]. On the other hand, the world’s knowledge is ever-growing, so the up-to-date knowledge is usually different from the one during LLMs’ pretraining [10]. Many such errors and emerging facts will arise sequentially in deployment, some of which have to be addressed timely and efficiently without waiting for retraining or finetuning [11, 12]. Also, retraining or finetuning is often too computationally expensive [13, 10], which is not sustainable for lifelong growing knowledge. Therefore, lifelong model editing [10] was proposed to remedy the continual knowledge updates and injections for LLMs in a cheap and timely manner.
An effective lifelong model editing approach should satisfy the following properties [14, 15, 11, 16, 17]: i) reliability, the model can remember both current and previous edits after sequential editing; ii) locality, model editing will not influence inherent pretrained knowledge which is irrelevant to the edited knowledge; iii) generalization, the model is not just merely memorizing the query-target pairs;
@@ 2.3.1 Side Memory Design
Side memory in FFN’s value matrix.
Each layer in a Transformer contains a multi-head self-attention (MHA) mechanism and a feed-forward network (FFN), where the FFN constitutes two-thirds of the model parameters [33]. The question of how Transformers retrieve and utilize stored knowledge remains unresolved [18, 34], yet past works [31, 33] have demonstrated that editing the weights of the FFN is consistently more effective for LLMs. The FFN typically consists of key-value linear matrices: $\mathbf{W}_{k},\mathbf{W}_{v}$, i.e., two multi-layer perceptron (MLP) layers. For the output of attention feature $\mathbf{f}$, the computation of the feed-forward network, omitting the bias terms, can be represented as:
@@
Locating side memory’s FFN layer.
Transformer LLMs have been widely demonstrated to encode “lower-level” information (e.g., parts of speech) in earlier layers while processing more advanced linguistic phenomena like anaphora and coreference in later layers [35, 36, 37]. Representations in later hidden layers propagate through residual connections without drastic changes [38, 18], enabling effective early exit in LLMs [39, 40]. Therefore, to minimize the side effects of editing and adjust advanced linguistic phenomena, we target mid-to-late layers (e.g. 27) for side memory. Further analysis of layer selection is provided in Section 3.3.
Routing between side memories and main memory.
Similar to the retrieval-based methods [10, 32], during inference, it is needed to decide whether the main memory or the side memory is used. If a given query is within the scope of previous edits, the side memory is used; otherwise, the main memory. Inspired by [11], we introduce a routing activation indicator, given an input $\mathbf{x}$, it is formulated:
@@ 2.3.2 Knowledge Sharding and Merging
Knowledge in random memory subspaces.
We edit the side memory $\mathbf{W}_{v^{\prime}}$. We divide $n$ edits into $k$ shards, copy the side memory for $k$ times, and generate $k$ random gradient mask with mask ratio $\rho$ for each copy of side memory. A random gradient mask $\mathbf{M}_{i}\in\{0,1\}^{|\mathbf{W}_{v^{\prime}}|},i\in[k]$ is a binary mask whose proportion of $1$ is $\rho$ [44]. For edit shard $i,i\in[k]$, we edit the knowledge into the subspace $\mathbf{M}_{i}$ as follows:
@@
Knowledge merging.
We merge the $k$ subspace pieces of side memory into one. Because we randomly generate the subspace masks, different random masks will have some overlapping elements and some disjoint elements, following the theorem below:
@@
Routing and retrieving among several side memories.
One single side memory has its limited knowledge capacity [43]. For the lifelong editing stream, we can produce several side memories and retrieve them via activation score routing. We compute different activation indicator scores of side memories and retrieve the top-1 during inference. This design is named WISE-Retrieve, which enables a more challenging lifelong editing scenario. For WISE with only one side memory, it is notated as WISE-Merge. For most of the experiments, we use WISE-Merge by default, and we compare WISE-Retrieve in Table 6 and Figure 6.
@@ 3.1 Experimental Settings and Evaluation Metrics
In the experiments, we compare the performance of different baselines and WISE in sequentially editing LLM models hundreds to thousands of times. In practice, we augment $\mathbf{x}_{e}$ by generating 10 random token sequences of length 10 using $f_{\Theta}$, enhancing editing generalization/adaptation to diverse contexts. We ensure that this augmentation with random tokens is applied across all baselines (See Appendix B.6, we ablate the contribution of Random Token).
@@
Datasets and Models.
We choose trending autoregressive LLM models LLaMA-2-7B [13], Mistral-7B [52], and GPT-J-6B [53, 54] for evaluation. The dataset details are in Table 3. Following [10], we evaluate WISE on the closed-book question-answering (QA) dataset ZsRE [46], and also evaluate its ability to correct Hallucination in SelfCheckGPT [48]. The Temporal dataset [50] is employed to test the out-of-distribution (OOD) generalization of editing. Since Temporal comprises emerging entities post-2019, we avoid using the latest LLMs in OOD experiments. Instead, we follow the original literature of the Temporal dataset [50] and adopt GPT-J-6B as the base model, which is pretrained on the Pile [51] with a cutoff in 2020. Implementation details and editing examples for each dataset and can be found in Appendix A.
Baselines.
The baselines include methods of continual learning and model editing. We compare WISE against direct fine-tuning FT-L with an additional KL divergence loss [18], and continual learning fine-tuning based on Elastic Weight Consolidation (FT-EWC) [20]. We also compare WISE to other model editors, including 1) GPT-style editors based on causal tracing: ROME [18], MEMIT [19], and MEMIT-MASS (a batch-editing version of MEMIT); 2) hypernetwork-based editors: MEND [31]; and 3) the latest memory-based editors: DEFER (inspired by SERAC [32] for inference routing) and GRACE [10]. Details on all comparisons are found in Appendix A.2.
Metrics.
Each edit example includes an edit descriptor (i.e., query) $\mathbf{x}_{e}$, its paraphrase prompts $\mathbf{x}_{e^{\prime}}$ (if available) for testing generalization, and an unrelated statement $\mathbf{x}_{\text{loc}}$ for testing locality. For the editing dataset $\mathcal{D}_{\text{edit}}=\{(\mathcal{X}_{e},\mathcal{Y}_{e})\}$ with $T$ edits, we evaluate the final post-edit model $f_{\Theta_{T}}$ after the $T$-th edit example $(\mathbf{x}_{T},\mathbf{y}_{T})$. We evaluate the model editor’s reliability and generalization using the metrics Rel. (a.k.a Edit Success Rate [10]) and Gen. (Generalization Success Rate [55]), while Loc. (Localization Success Rate [55]), defined as the post-edit model should not change the output of the irrelevant examples $\mathbf{x}_{\text{loc}}$, assesses specificity. We report these metrics and their mean scores, which are formally defined as:
@@ 3.2 Main Results
Competitive Performance of WISE.
The competitive performance of WISE is evident in Table 2 and 4, which compare its results with eight baselines on the QA (ZsRE) and Hallucination (SelfCheckGPT) settings. In general, we observe the followings: ❶ WISE outperforms existing methods on multiple tasks after long editing sequences; ❷ direct editing of long-term memory (ROME, MEMIT, etc.) creates conflicts with prior pretraining knowledge, resulting in poor locality; and ❸ retrieving working memory and modifying activations (GRACE, DEFER, etc) struggle to generalize to diverse queries.
@@
Out-of-Distribution Evaluation.
Ideally, model editing needs to generalize distributionally from formulaic editing examples to natural texts [50], where the distributional shift involves complexity rather than conventional domain shift [56]. Following [50], we evaluate the OOD generalization of editing methods on emerging entities using the temporal updating dataset, Temporal. Editing examples and evaluation metrics are provided in Appendix A.1. As shown in Table 5, WISE effectively handles out-of-distribution generalization tasks (achieving the best OOD Gen. and overall performance). DEFER delivers mediocre performance on OOD Gen. due to the limited capacity of the auxiliary model[14]. During the fine-tuning phase, GRACE and MEMIT focus on the representation $v*$ of a single input token after $\mathbf{W}_{v}$ (GRACE: last token, MEMIT: last subject token). However, regarding $v*$ the editing carrier encounters two problems: 1) the training objective is not aligned with the pretraining phase, and 2) the single representation limits the search scope of gradient descent, making it difficult to handle OOD generalization. WISE, on the other hand, avoids these challenges.
@@ 3.3 Further Analysis
Visualization of WISE’s Routing Activation.
To demonstrate the effectiveness of memory routing, we record the activation values $\Delta_{\text{act}}(\mathbf{x})$ of 1000 (QA, ZsRE)/600 (Halluc.) queries during the inference stage via knowledge merging into a single side memory. As shown in Figure 3, the purple horizontal line represents the activation threshold $\epsilon$ recorded during the editing phase. Almost all unrelated queries show low activations with values less than 10 in ZsRE and less than 20 in Halluc.; meanwhile, WISE accurately routes the editing prompt and unseen paraphrases into the side memory. This ensures editing locality and prevents excessive shifts from the pre-training distribution during lifelong editing.
@@
Localization Analysis of WISE’s Side Memory.
To validate the benefits of editing mid-to-late layers, we select decoder layers from early, intermediate, mid-to-late, and late stages. As shown in Figure 5, the ablation results reveal that editing critical layers like the early and final layers (0, 1, 31) is ineffective, even resulting in a very low Loc. value of 0.096, which indicates a failure to recognize the editing scope. This may occur because the early layers represent fundamental grammatical information, and the final layer directly controls the decoding procedure, leading to poor editing of advanced language functions. Editing in the intermediate layers is suboptimal but still shows a markable improvement compared to early layers, possibly because intermediate layers start to integrate basic grammatical information with more complex semantic data. Notably, the mid-to-late layers demonstrate exceptional editing performance; for instance, selecting layer 26 results in an 80% success rate and generalization while maintaining 100% locality. This empirically supports our claim in Section 2.3.1 that the redundant mid-to-late layers [39] are ideal side memory layers and confirms the hierarchical nature of information processing in Transformer LLMs [57, 58].
Analysis of $\rho$ and $k$ for WISE.
We analyze the important hyperparameters of WISE: the mask ratio $\rho$ and the number of subspaces $k$ in Figure 5. On the left figure, for $k=2$, the best $\rho$ is 0.2, satisfying $k*\rho=0.4<1$, which implies the effectiveness of our subspace design that higher knowledge density will cause better generalization. When scaling $k$, we observe an increasing demand of $\rho$. From Theorem 2.1, the probability of subspace overlap is $\rho^{k}$, and we hypothesize that this overlap is important as an anchor for model merging. Interestingly, from the right figure, it can be observed that the optimal cases always have the $\rho^{k}$ closest to 0.03. This shows an inherent tradeoff between merge anchor and merge conflicts, and the subspace overlaps around 0.03 are optimal for the best performances. Such experiments indicate that 20% FFN parameters can accommodate at least 500 edited samples. When "mask memory exhaustion" occurs, we can allocate new mask parameters to store new knowledge. Using retrieve when knowledge isn’t full and merging as needed to save memory, achieves true lifelong model editing.
@@
Scale Up to 3K of Edits.
We scale the number of continual edits to 3K in Table 6. We compare WISE-Merge, keeping one side memory by multi-time merging, and WISE-Retrieve, keeping several side memories by routing and retrieving among different side memories. For WISE-Retrieve, we show an upper bound “oracle”, which always identifies the correct routing path. We observe that the WISE series maintains high scalability, consistently outperforming the strongest baselines including MEMIT-MASS and GRACE. WISE-Retrieve based on top-1 activation retrieval demonstrates the best results in 3K edits, showing the effectiveness of well-organized memory subspaces and routing strategies during editing. We note that the “oracle” exhibits marginal performance decline when scaling the edits from 2K to 3K, yet it demonstrates remarkable performance across all metrics. This underscores the potential of WISE to handle extremely long continual edits, contingent upon substantial improvement in the retrieval of side memories. Additionally, an appropriate replay of edits can further improve retrieval accuracy, as detailed in Appendix B.3.
Table 7: Ablation study of Router (compared with Table 2). LlaMA.
$\text{WISE}_{\text{w.o.}\,L_{a}}$ Rel. Gen. Loc. Avg.
$T=1$ 1.00 0.96 0.93 -0.07 0.96 -0.01
$T=10$ 0.93 0.90 0.88 -0.12 0.90 -0.04
$T=100$ 0.92 0.85 0.81 -0.19 0.86 -0.04
$T=1000$ 0.84 0.79 0.72 -0.28 0.78 -0.05
Contribution of Router designs in WISE.
Without the router strategy, all inputs either pass solely through the main or side memory. To further validate its effectiveness, we conduct additional ablations with $L_{a}$. WISE’s performance on ZsRE is shown in Table 7. We observe the expected decrease in Loc. w.o. $L_{a}$, such as dropping from 1.00 to 0.72 at T=1000, reveals the router’s effectiveness in identifying editing scopes, minimizing side effects, and retaining a substantial amount of pre-training knowledge.
@@
Inference Time Analysis of WISE.
Figure 6 shows the inference time of a single instance for LLaMA after $t\in[0,3000]$ editing steps, measured across 10 trials of each setting. Consistent with our expectations, we find that WISE-Merge incurs a constant inference delay (about 3%) as the editing stream expands. WISE-Retrieve, due to the introduction of retrieval routing, shows an increase in inference time as the number of edits increases, with a time cost increment of about 7% after 3K edits. Knowledge merging ensures that WISE-Merge only brings constant additional costs (0.64% extra parameters and 4% extra GPU VRAM, as detailed in Appendix B.7), contrasting with past memory-based works that continuously demand more available memory [10, 32].